a students is taking five course and is facing the crunch of final exams.she estimates that she has 40 hours available to study .if xj=the number of hours allocated to study for coursej.state the equation whose solution set specifies all possible allocation of time among the five courses which will exhaust the 40 hours
1
Expert's answer
2020-05-11T20:01:31-0400

Student has five subjects and 40 hrs available.

So,"x_1+x_2+x_3+x_4+x_5=40"

"0\leq x_1\leq40, \,0\leq x_2\leq40,\\ 0 \leq x_3\leq40,0\leq x_4\leq40,0 \leq x_5\leq40,"

where "x_1,x_2,x_3,x_4,x_5" are real numbers.

This is the equation whose solution set specifies all possible allocation of time among the five courses which will exhaust the 40 hours.